OpenStudy (jhannybean):

Just working on some practice problems. \(\int\int_R (4x+8y)dA\) where \(R\) is a parallelogram with vertices (-1,3),(1,-3) and (3,-1), (1,5); \(x=\frac{1}{4}(u+v)~,~ y=\frac{1}{4}(v-3u)\)

OpenStudy (jhannybean):

Hm.. 

ganeshie8 (ganeshie8):

you want to change parallelogram into a rectangle ?

OpenStudy (jhannybean):

Yeah, I guess.

OpenStudy (jhannybean):

I'm not sure where to begin exactly...can I get a hint? lol

OpenStudy (jhannybean):

Oh wait, do we even need to? We can just use these two equations to get the jacobian, right?

ganeshie8 (ganeshie8):

yeah looks you're given the transformation already

OpenStudy (jhannybean):

How do you write up matrices...

OpenStudy (jhannybean):

@ParthKohli transformations from cartesian to polar

ganeshie8 (ganeshie8):

``` \begin{vmatrix} a&b\\ c&d \end{vmatrix} ``` produces \[\begin{vmatrix} a&b\\ c&d \end{vmatrix}\]

OpenStudy (jhannybean):

\[\left|\frac{\partial(x,y)}{\partial(u,v)}\right|=\left[\begin{matrix}\frac{1}{4} & \frac{1}{4} \\ \frac{1}{4} & \frac{1}{4}\end{matrix}\right] \]

OpenStudy (jhannybean):

Why am I not getting this..

ganeshie8 (ganeshie8):